# When in Doubt, Consult: Expert Debate for Sexism Detection via Confidence-Based Routing

### v2

@@ 4 Experiments
This section describes the datasets, evaluation metrics, experimental setup, and results used to assess the proposed framework.
@@ 4.1 Experimental Setup
Datasets. We evaluate our methods on two widely used benchmark corpora for sexism detection: the EXIST 2025 Tweets dataset exist25 and the EDOS dataset edos23 (detailed descriptions in Appendix A.1). The EXIST dataset is selected for its difficulty, as it incorporates annually updated tweets, capturing emerging sexist language, evolving slang, and annotation disagreements that mirror genuine ambiguities in human judgment. The EDOS dataset offers a complementary perspective with high quality fine-grained annotations at multiple levels of granularity, organized across three hierarchical classification tasks.
Evaluation Metrics. Following the official setup, we use the Evaluate ALL 2.0 tool evall2025 for scoring under the EXIST 2025 Task 1.1 (Hard-Hard) configuration. The evaluation metrics include: (i) Inter-Consistency Measure (ICM) amigo2022evaluating, which quantifies agreement between predicted and reference annotations by combining intra-system consistency and penalizing overlap; (ii) ICM-Norm, a normalized version of ICM scaled to $[0,1]$ for comparability; and the (iii) F1 score, the harmonic mean of precision and recall. As for the EDOS dataset, we follow the official evaluation protocol and report results using the Macro-F1 score, the unweighted mean of the F1-scores computed independently for each class across Task A, B, and C (Appendix A.1).
Baselines. We evaluate our proposed framework against both fine-tuned and zero-shot LLM baselines. These focus on state-of-the-art approaches for sexism detection; while an exhaustive exploration of all available LLMs is beyond the scope of this work. As specialized models, we employ LLaMA-3.2-3B, adapted via LoRA fine-tuning (hu2021lora) on both the EDOS and EXIST training datasets. Full task-specific training information is available in Appendix A.2. To evaluate instruction-following and reasoning abilities in a zero-shot setting, we test Qwen2.5-72B-Instruct qwen2025qwen25technicalreport, LLaMA-3.3-70B-Instruct grattafiori2024llama, GPT-5.2 openai2025gpt5, and Cogito-70B deepcogito_cogito_v2 using simple classification prompts (Appendix A.3). Pre-trained checkpoints were obtained from HuggingFace huggingface. All LLM experiments were executed locally through LangChain langchain and Ollama ollama on a system equipped with three Nvidia A100 GPUs.
@@ 4.2 Results
In this section, we present the experimental results of our framework and baselines on the EXIST 2025 and EDOS benchmarks.
@@
EXIST Dataset. Results for the EXIST 2025 Task 1.1 test set, binary classification, are reported in Table 1. The first group of models comprises LLMs evaluated in a zero-shot setting. Among these, Cogito-70B (reasoning mode) achieves the highest F1_YES (0.7143), outperforming Qwen2.5-72B-Instruct (0.7112), standard Cogito-70B (0.7109), and state-of-the-art GPT-5.2 (0.7043). Overall, all zero-shot models with basic prompts underperform compared to leaderboard approaches, reflecting the limitations of out-of-the-box LLMs for specialized tasks.
The second group reports official leaderboard systems submitted to EXIST 2025 Task 1.1. Here, LLaMA-3.1-8B-Instruct tian2025marioexist2025simple leads in both ICM (0.6774) and F1_YES (0.8167). Transformer fusion architectures and BERT-based systems round out the group, showing the persistence of competitive results of earlier architectures.
@@
The final section of Table 1 details the results of our routing framework. We build upon the LLaMA-3.2-3B speciliazed model which achieves performance on par with lower-ranking leaderboard models (F1_YES: 0.7596). Applying targeted optimizations in $\mathcal{C}_{2}$ (multi-dataset training, class-balanced loss, threshold calibration) yields substantial improvements: 0.8089 F1_YES and 0.8257 macro-F1, with results that already surpass many leaderboard submissions. Further gains are realized through confidence-aware dynamic routing. Configuration $\mathcal{C}_{3}$ integrates Qwen2.5-72B-Instruct for persona-based debate, reaching 0.8191 F1_YES, while $\mathcal{C}_{4}$, leveraging LLaMA-3.3-70B for personas and Cogito-70B as judge, achieves the highest scores across all reported metrics. This configuration outperforms the best leaderboard system by 0.66 points on F1_YES and establishes a new benchmark for the task.
EDOS Dataset. We evaluate the results on the EDOS dataset, reporting the macro-F1 across three classification tasks of increasing granularity (Table 2). Results are presented for zero-shot LLMs, official leaderboard systems, models employing data augmentation and ensembling, and our selective routing framework.
@@
The first group reports the results of LLMs evaluated in a zero-shot setting. Performance remains limited across all tasks when using basic prompts. GPT-5.2 achieves the highest performance across all tasks (Task A: 0.7306, Task B: 0.4543, Task C: 0.3381), followed by LLaMA-3.3-70B-Instruct (0.6421). Cogito-70B in reasoning mode consistently outperforms its standard variant (+2.26 pp in Task A). However, all zero-shot models struggle with fine-grained classification (Task C < 0.34).
Among official leaderboard systems from SemEval 2023 Task 10.A, ensemble methods based on DeBERTa and RoBERTa architectures set a strong baseline, reflecting the effectiveness of large pre-trained transformer models with supervised fine-tuning.
The next group encompasses systems employing data augmentation, model ensembling, or both. Here, CSE khan2025explaining and DDA khan2025explaining achieve notable gains: CSE reaches 0.8819 on Task A and 0.5639 on Task C, while DDA records the highest reported macro-F1 on Task C (0.6018) and competitive results on other tasks.
@@
Our approach in the base configuration, $\mathcal{C}_{1}$ (fine-tuned LLaMA-3.2-3B), achieves competitive results for binary and multi-class tasks, but falls short on fine-grained Task C (0.3717). Introducing targeted training optimizations in $\mathcal{C}_{2}$, including multi-dataset training, class-balanced loss, and threshold calibration, yields marked improvements, reaching 0.8986 on Task A, 0.6279 on Task B, and 0.4189 on Task C. Finally, confidence-aware escalation to the CEJ module provides substantial further gains. Configuration $\mathcal{C}_{3}$, which employs Qwen2.5-72B-Instruct for persona-based debate, achieves 0.9195 on Task A and 0.7324 on Task B. The strongest results are obtained by $\mathcal{C}_{4}$, where LLaMA-3.3-70B is used for the persona debate and Cogito-70B as the judge, achieving 0.9214 on Task A and 0.7421 on Task B. Both configurations outperform all previous systems on binary and multi-class classification. On Task C, both routing variants remain competitive (0.5904 and 0.5842), though they do not surpass the best-performing data augmentation method (DDA). This gap can be attributed to the cascading effect of baseline limitations. Despite employing task-specific optimization techniques, ($\mathcal{C}_{2}$) exhibits progressive macro-F1 score degradation as class granularity increases (0.8986 $\rightarrow$ 0.6279 $\rightarrow$ 0.4189 across Tasks A, B, and C), indicating that these optimization techniques only partially mitigate the challenge of learning 11 fine-grained categories with severe class imbalances. Since CEJ routes instances based on $\mathcal{C}_{2}$’s confidence scores, this weaker starting point propagates through the pipeline. In contrast, DDA addresses data scarcity directly by synthesizing training examples that populate underrepresented categories, providing a stronger foundation that inference-level refinement alone cannot replicate.

### v3

@@ 4 Experiments
We proceed describing the datasets, evaluation metrics, and results used to assess the proposed framework.
@@ 4.1 Experimental Setup
Datasets. We evaluate our methods on two widely used benchmarks for sexism detection: the EXIST 2025 Tweets dataset Plaza et al. (2025) and the EDOS dataset Kirk et al. (2023). The EXIST dataset incorporates annually updated social media posts, capturing emerging sexist language, evolving slang, and annotator disagreements that mirror ambiguities in human judgment. The EDOS dataset offers a complementary perspective with high quality fine-grained annotations at multiple levels of granularity, organized across three hierarchical classification tasks (see Appendix A.1).
Evaluation Metrics. Aligned with the official setup, we use the Evaluate ALL 2.0 tool de Educación a Distancia (UNED) (2025) for scoring under the EXIST 2025 Task 1.1 (Hard-Hard) configuration. The evaluation metrics include: (i) Inter-Consistency Measure (ICM) Amigo and Delgado (2022), which quantifies agreement between predicted and reference annotations by combining intra-system consistency and penalizing overlap; (ii) ICM-Norm, a normalized version of ICM scaled to $[0,1]$; and the (iii) F1 score.As for the EDOS dataset, we follow the public evaluation protocol, reporting results for the Macro-F1 score, the unweighted mean of the F1-scores computed independently for each class across task A, B, and C (Appendix A.1).
Baselines. We evaluate our proposed framework against both zero-shot and fine-tuned LLM baselines. It worth noting that we focus on state-of-the-art approaches for sexism detection rather than aiming at an exhaustive exploration of all available LLMs, which would be beyond the scope of this work. As specialized models, we employ LLaMA-3.2-3B, adapted via LoRA fine-tuning (Hu et al., 2021) on both the EDOS and EXIST training datasets. Full task-specific training information is available in Appendix A.2. To evaluate instruction-following and reasoning abilities in a zero-shot setting, we test Qwen2.5-72B-Instruct Qwen et al. (2025), LLaMA-3.3-70B-Instruct Grattafiori et al. (2024), GPT-5.2 OpenAI (2025), and Cogito-70B Cogito (2025) using consistent classification prompts (Appendix A.3). Pre-trained checkpoints were obtained from HuggingFace Wolf et al. (2020). All experiments were conducted locally through LangChain Chase (2022) and Ollama Team (2023), on a system equipped with 3 Nvidia A100 GPUs.
@@ 4.2 Results
EXIST Dataset. Results for the binary classification task 1.1 of EXIST 2025 are reported in Table 1. The first group of models comprises LLMs evaluated in a zero-shot setting. Among these, Cogito-70B (reasoning mode) achieves the highest F1_YES (0.7143), outperforming by little Qwen2.5-72B-Instruct (0.7112), standard Cogito-70B (0.7109), and GPT-5.2 (0.7043). Overall, all zero-shot models with simple prompts still underperform smaller yet specialised approaches, reflecting still the limitations of LLMs prompted with simple requests. The second group reports public leaderboard systems submitted to EXIST 2025 Task 1.1. Here, LLaMA-3.1-8B-Instruct Lin Tian and Rizoiu (2025) leads in both ICM and F1_YES. Systems based BERT-variants still shows competitive results compare to larger architectures with simple prompts.
@@
The bottom of the Table 1 details the results of our routing framework. The initial specialised model, LLaMA-3.2-3B, achieves performance on par with lower-ranking leaderboard models (F1_YES: 0.7596). Applying simple but targeted regularization, indicated by $\mathcal{C}_{2}$ (multi-dataset training, class-balanced loss, threshold calibration), yields already substantial improvements: 0.8089 F1_YES and 0.8257 macro-F1, with outperforming several leaderboard submissions. Further gains are realized through the dynamic routing based on confidence calibrations. Configuration $\mathcal{C}_{3}$ integrates Qwen2.5-72B-Instruct for persona-based debate, reaching 0.8191 F1_YES, while $\mathcal{C}_{4}$, leveraging LLaMA-3.3-70B for personas and Cogito-70B as judge, achieves the highest scores across all reported metrics. This configuration, achievement state-of-the-art peformance on the task, demonstrates the importance of controlling different LLMs to leverage wider capabilities for the task.
EDOS Dataset. For the EDOS dataset, we report in Table 2 the Macro-F1 across the three classification tasks of increasing granularity.
@@
The first group reports the results of LLMs evaluated in a zero-shot setting. Performance remains limited across all tasks when using simple prompts, with GPT-5.2 achiving the highest performance across all tasks, followed by the open-weight LLaMA-3.3-70B-Instruct. Cogito-70B in reasoning mode consistently outperforms its standard variant. However, all zero-shot models particularly struggle with fine-grained classification (Task C). Among official leaderboard systems from SemEval 2023 Task 10.A, ensemble methods based on DeBERTa and RoBERTa architectures are also in this case strong baselines. The next group reports systems employing data augmentation, model ensembling, or both. CSE Khan et al. (2025a) and DDA Khan et al. (2025a) achieve notable gains: CSE reaches 0.8819 on Task A and 0.5639 on Task C, motivating the data expansion and leveraging perspectives from multiple models.
@@
Our approach in the base configuration, $\mathcal{C}_{1}$, i.e., fine-tuned LLaMA-3.2-3B, achieves competitive results for binary and multi-class tasks despite its simplicity, but falls short on fine-grained tasks. Introducing targeted training regularization also here, $\mathcal{C}_{2}$, yields again marked improvements, outperforming the zero-shot baselines and most of the specilised systems. Finally, confidence-aware escalation to the CEJ module provides substantial gains. Configuration $\mathcal{C}_{3}$, which employs Qwen2.5-72B-Instruct for persona-based debate, achieves 0.9195 on Task A and 0.7324 on Task B, with the strongest results are obtained by $\mathcal{C}_{4}$, where LLaMA-3.3-70B is used for the persona debate and Cogito-70B as the judge. Both configurations outperform all previous systems on binary and multi-class classification on task A and B. On Task C, both routing variants remain competitive, though they struggle compare to the DDA data augmentation method. This gap can be attributed to the cascading effect of baseline limitations: despite employing task-specific optimization techniques, ($\mathcal{C}_{2}$) exhibits progressive macro-F1 score degradation as class granularity increases (0.8986 $\rightarrow$ 0.6279 $\rightarrow$ 0.4189 across Tasks A, B, and C), indicating that these optimization techniques only partially mitigate the challenge of learning 11 fine-grained categories with severe class imbalances. In contrast, DDA addresses data scarcity directly by synthesizing training examples that populate underrepresented categories, providing a stronger foundation that inference-level refinement alone cannot replicate.